The odds against a horse winning a race are 6 to 5. What is the probability of the horse winning the race?
Answer by Nate(3500)

The odds against a horse winning a race are 6 to 5.
Odds are taken as: chances of winning: chances of losing
Probability: chances of winning / total chances
Prob: 6 / (6 + 5) = 6 / 11 = about 54.5%